Answer the following questions on this sheet or in your laboratory notebook.


The reaction for the formation of diamminesilver ion is as follows: Ag + (aq) + 2 NH3 (aq)  Ag(NH3)2 +


(2 points). 1. Write the equilibrium constant expression for the reaction


2. An experiment was carried out to determine the value of the equilibrium constant, Kc for the reaction.  Total moles of Ag+ present= 3.6 x 10 -3 moles  Total moles of NH3 present = 6.9 x 10-3 moles  Measured concentration of Ag(NH3)2 + at equilibrium = 3.4 x 10-2 M  Total solution volume = 100 mL

a. (2 points). Calculate the equilibrium concentration of Ag+ (uncomplexed)







b. (3 points). Calculate the equilibrium concentration of NH3 (uncomplexed)







c. (3 points). Calculate the value of the equilibrium constant
Answer #1

1. Write the equilibrium constant expression for the reaction

K = [Products]/[Reactants] = [Ag(NH3)2+]/[Ag+][NH3]2

2.

                              Ag+(aq) + 2NH3(aq)         \rightleftharpoons            Ag(NH3)2+

                               A                  B                                   C

Ag+(aq)

2NH3(aq)                     

Ag(NH3)2+

Initial

CAo

CBo

CCo

Equilibrium

CA = CAo(1-x)

CB = CBo- 2AoX

CC = CCo + CAoX

Given

CC = CAoX = 3.4 x 10-2 M

But, CAo = (3.6 x 10-3 moles)/ 0.1 L = 0.036 M

Therefore, X = (3.4 x 10-2)/(3.6 x 10-2) = 0.944

a. Calculate the equilibrium concentration of Ag+ (uncomplexed)

Equilibrium concentration of Ag+ = CA = CAo (1-X) = (3.6 x 10-2) (1- 0.944) = 0.002 M

b. Calculate the equilibrium concentration of NH3 (uncomplexed)

Equilibrium concentration of NH3 = CB = CBo -2 CAoX

But, CBo = (6.9 x 10-3 moles)/ 0.1 L = 0.069 M

CB = CBo -2 CAoX = 0.069 M – (2 x 0.036 x 0.944) = 0.001 M

c. Calculate the value of the equilibrium constant

K = [Products]/[Reactants] = [Ag(NH3)2+]/[Ag+][NH3]2

                                                = (3.4 x 10-2)/[(2 x 10-3)(1 x 10-3)2] = 1.69 x 10-7